OpenStudy (anonymous):

Determine points when the function is concave up and down and find inflection points F(x)=-3x^3+-4x^2+1x-11

OpenStudy (zpupster):

www.mathportal.org/calculators/polynomials-solvers/polynomial-graphing-calculator.php?val1=3x^3+plus4x^2plusx-11&ch1=xi&ch5=inf&ch6=con 1st der 9x2+8x+1 2nd der 18x + 8 set 0 = 18x+8 x=-4/9 back in equation y= - 10.91 that is the inflection point you determine concave up or down by picking numbers larger and smaller and plugging them in and seeing what the graph is doing.